OpenStudy (anonymous):

How is the graph of y=(x+9)^2 transformed form the graph of y=x^2.

jimthompson5910 (jim_thompson5910):

y=(x+9)^2 is the same as y=(x-(-9))^2 if you compare y=(x-(-9))^2 to y=a(x-h)^2 +k, you'll find that a = 1, h = -9 and k = 0

jimthompson5910 (jim_thompson5910):

what does that value of h tell you?

OpenStudy (anonymous):

How many units your going to move?

jimthompson5910 (jim_thompson5910):

yes you're on the right track

jimthompson5910 (jim_thompson5910):

does h tell you how to move up and down? or does h tell you how to move left and right?

OpenStudy (anonymous):

move up and down? right?

jimthompson5910 (jim_thompson5910):

no, that's actually what k does

OpenStudy (anonymous):

ohh so it tells which way to move either left or right?

jimthompson5910 (jim_thompson5910):

yes, if h is negative, are we going left or right?

OpenStudy (anonymous):

we would be going left

jimthompson5910 (jim_thompson5910):

how many units to the left

OpenStudy (anonymous):

9

jimthompson5910 (jim_thompson5910):

correct to go from y = x^2 to y = (x+9)^2, you shift the original parent graph 9 units to the left

OpenStudy (anonymous):

ohh ok i see now